OpenStudy (anonymous):

The perimeter of an isosceles right triangle is 16+16√2, what is the length of the hypotenuse?

OpenStudy (anonymous):

16 sqrt(2).

OpenStudy (anonymous):

Let x be the leg length. Then the hypotenuse is x sqrt(2), and the perimeter is 2x + x sqrt(2) = x(2 + sqrt(2)) = 16(1 + sqrt(2)). Solve for x.